If β is such that sin2 β is not equals to zero, then show that the expression [x 2 +2x cos2α+1] / [x 2 +2x cosβ +1 ] for x belonging to R always lie between cos 2 α / cos 2 β and sin 2 α/sin 2 β.

If   β is such that sin2 β is not equals to zero, then show that the expression


[x2+2x cos2α+1]   / [x2+2x cosβ +1 ] for x belonging to R always lie between cos2α / cos2β  and sin2α/sin2β.


 


 

Grade:12

2 Answers

Ramesh V
14 years ago

Lets take y=[x2+2x cos2α+1]/[x2+2x cos2β +1] for x belonging to R

so, its [x2+2x cos2α+1] - y*[x2+2x cos2β +1] = 0

       (1-y)x2+2x(cos2α-cos2β)+(1-y) = 0

       the equation has real roots, when discriminant os positive

i.e., (cos2α-cos2β)2 - (1-y)2 > 0

      [(cos2α-cos2β)+(1-y)].[(cos2α-cos2β)-(1-y)] > 0

       [ y(1+cos2β)-(1+cos2α) ].[ y(1-cos2β)-(1-cos2α) ] < 0

           (1+cos2α)/(1+cos2β)  <  y  <  (1-cos2α)/(1-cos2β)

 so, y always lie between cos2α/cos2β < y < sin2α/sin2β.

--

Please feel free to post as many doubts on our disucssion forum as you can. If you find any question difficult to understand - post it
here and we will get you the answer and detailed solution very quickly.We are all IITians and here to help you in your IIT JEE
preparation. All the best.

Regards,
Ramesh
IIT Kgp - 05 batch



mycroft holmes
14 years ago

The given expression can be written as 1 + 2(cos 2A - cos 2B)/(x + 1/x + 2 cos 2B)

 

Since x + 1/x >= 2 or <= -2, we have  1/(cos 2B -1) <= 1/(x+1/x + 2 cos 2B) <= 1/2(1+cos 2B)

 

Hence 2(cos 2A - cos 2B)/(x + 1/x + 2 cos 2B) lies between 2(cos 2A - cos 2B)/2(1+cos 2B) and 2(cos 2A - cos 2B)/2(cos 2B-1)

 

so that 1 + 2(cos 2A - cos 2B)/(x + 1/x + 2 cos 2B) lies between 2(1+cos 2A)/2(1+cos 2B) and 2(1-cos 2A)/2(1-cos 2B) i.e.

 

between cos2A/cos2B and sin2A/sin2B